OpenStudy (anonymous):

Find the angle between vectors u=<-1,5> and v=<3,-2>. Help!!! Please!! Will give MEDAL!!

OpenStudy (anonymous):

angle between two vectors is the dot product, isnt' it? http://www.mathsisfun.com/algebra/vectors-dot-product.html

OpenStudy (anonymous):

So 4.9??????

OpenStudy (anonymous):

I got 135.1 deg what did you get for your dot product of the two vectors? and what did you get for |a| * |b|?

OpenStudy (anonymous):

I got a=4 an B=2.2 I know that wrong so how did you get A and B?

OpenStudy (anonymous):

so its like this: u (dot) v = |u| * |v| cos(angle between) u (dot) v = x_u * x_v + y_u * y_v if vector u=<-1,5> if vector v=<3,-2> then u (dot) v = = -1*3 + 5*2 =?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

ya now do the cross product of u and v

OpenStudy (anonymous):

|u| x |v|

OpenStudy (anonymous):

like <-1,5>*<3,-2>???<3,-10>

OpenStudy (anonymous):

to do a cross product, you find the magnitutde of both vectors and multiply them together i'll find the magnitude of vetor u, sqr(-1^1 + 5^2) = sqr(1+25) = 5.099 whats the magnitude of vector v?

OpenStudy (anonymous):

2.236???

OpenStudy (anonymous):

sqrt(3^2 + -2^2) =sqrt(9+4) =3.6

OpenStudy (anonymous):

u (dot) v = |u| * |v| cos(angle between) -13 = 5.09*3.6 * cos(angle between)

OpenStudy (anonymous):

So it its the absolute value of it??

OpenStudy (anonymous):

ohhh, okok when we are dealing with vectors and we have the absolute value things like this: |v| it means to find the magnitude, aka lenght of the vector remember triangles and when we did pythagorean theorem? a^2 + b^2 = c^2? this is all finding the magnitude is, we are given the lenghts for side 'a' and side 'b' as givn in the vector 

OpenStudy (anonymous):

Okay I see how you got -13=5.09*3.6*cos now.. So what would you do after that???

OpenStudy (anonymous):

now we isolate cos() -13=5.09*3.6*cos -13/(5.09*3.6) = cos(angel between) then we do arc cosine, or inverse cosine cos^-1(-13/(5.09*3.6) ) = angle between

OpenStudy (anonymous):

135.1

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thank You !! Lots!!! you don't know but this has been very helpful!